What is the specific heat of a substance if 19.5 kJ are required to raise the temperature of 205 g of the substance from 56.5°C to 197.0°C?
1
Expert's answer
2012-10-08T10:24:09-0400
Q=m*c*(t2-t1)
c=Q/(m*(t2-t1))=19.5 kJ / (0.205 kg *(197-56.5)K) = 0.677
kJ/(kg*K)